OpenStudy (noseboy908):

Pre-Calculus: Find a formula for the nth term: Given that a(4)=-12, & a(5)=-6

OpenStudy (noseboy908):

@Johnbc If you can help, please take a look.

OpenStudy (anonymous):

We have the 4th and 5th term so we can easily find the difference

OpenStudy (anonymous):

After we have the difference you can use the same method we used last problem to find the A1 term

OpenStudy (noseboy908):

The difference would be 1/2

OpenStudy (anonymous):

\[A_n=A_1 + (n-1)d\]

OpenStudy (anonymous):

Yes it did go half or it also went +6

OpenStudy (noseboy908):

I forgot to mention, but it's geometric. That's also given.

OpenStudy (anonymous):

Geometric Sequence\[A_n= A_1 \times r^{n-1}\]

OpenStudy (noseboy908):

I believe I found the answer, I'll post the steps I used.

OpenStudy (anonymous):

 This should be similar to your steps?

OpenStudy (noseboy908):

Root formula: t(n)=t(1) x r^(n-1) -12=t(1) x (1/2)^(4-1) -12=t(1)(.125) -96=t(1) So, the formula would be "t(n)=-96+(1/2)^(n-1)

OpenStudy (anonymous):

Correct!

OpenStudy (noseboy908):

Alright, thanks again! I may be back with more momentarily.